# Semitoric Families on Pentagon Spaces

### v1

@@ 4. Non-degeneracy of non-transition points
$W_{1}(\widetilde{W}_{i}H_{t}^{\textrm{red},c})(\rho)=-tg_{3}^{i}(\rho)\frac{b_{4}}{b_{3}}+(1-t)g_{5}^{i}(\rho)\frac{b_{4}}{b_{5}},$
which evaluates to $\frac{b_{4}}{b_{3}}t-\frac{b_{4}}{b_{5}}(1-t)$ when $i=1$ and evaluates to 0 when $i=2$.
Similarly, one can compute
$W_{2}(\widetilde{W}_{2}H_{t}^{\textrm{red},c})(\rho)=\frac{a(a_{3}b_{5}-a_{5}b_{3})^{2}}{c}-t\frac{a_{3}^{2}+b_{3}^{2}}{b_{3}}b_{4}b_{5}^{2}-(1-t)\frac{a_{5}^{2}+b_{5}^{2}}{b_{5}}b_{4}b_{3}^{2}.$
Thus, the matrix representation of the Hessian of $H_{t}^{\textrm{red},c}$ at $Q$ under the basis $W_{1}],[W_{2}]$ is
$\begin{pmatrix}-t\frac{b_{4}}{b_{3}}-(1-t)\frac{b_{4}}{b_{5}}&0\\ 0&\frac{a}{c}(a_{3}b_{5}-a_{5}b_{3})^{2}-\frac{t}{b_{3}}(a_{3}^{2}+b_{3}^{2})b_{4}b_{5}^{2}-\frac{1-t}{b_{5}}(a_{5}^{2}+b_{5}^{2})b_{3}^{2}b_{4}\end{pmatrix}$
We show that this matrix has positive determinant. Set $X=a_{4}b_{3}-a_{3}b_{4},\ Y=a_{4}b_{5}-a_{5}b_{4}$, then $X+Y=a_{4}(b_{3}+b_{5})-(a_{3}+a_{5})b_{4}=-a_{4}b_{4}+(c+a_{4})b_{4}=cb_{4}$. Since $c\neq 0$, we can write $b_{4}=\frac{X+Y}{c}$. By Equation (*4.1), $ab_{3}=tX,ab_{5}=(1-t)Y$, so we have $X=\frac{ab_{3}}{t},Y=\frac{ab_{5}}{1-t}$. Hence,
@@
Let $K=\frac{Y}{X}=\frac{tb_{5}}{(1-t)b_{3}}$, then $A=-\frac{a}{cK}(K+1)^{2}$. Similarly, one can compute
@@
By item (2) of Lemma 4.1, $K\neq-1$, so $A,B$ are both nonzero. Hence, $\det(\operatorname{Hess}(H_{t}^{\textrm{red},c}))=AB$ has the same sign as $\dfrac{a^{2}}{c^{2}K^{2}}>0$. We conclude that $Q$ is an elliptic singularity (in the Morse sense).
Case 2: when $b_{4}=0$, then $a_{4}\neq 0$, so we choose the charts with coordinates $(y_{0},z_{0},x_{3},z_{3},y_{4},z_{4},x_{5},z_{5})$. Notice that by (1) of Lemma 4.1, this could only happen when $t=\dfrac{1}{2}$. A similar computation shows that under the basis $W_{1}=\frac{\partial}{\partial z_{0}}-\frac{\partial}{\partial z_{3}}$ and $W_{2}=(a_{5}b_{3}-a_{3}b_{5})\frac{\partial}{\partial y_{0}}-b_{3}b_{5}\frac{\partial}{\partial x_{3}}+b_{3}b_{5}\frac{\partial}{\partial x_{5}}$, the matrix representation of the Hessian of $H_{t}^{\textrm{red},c}$ is given by
$\begin{pmatrix}\dfrac{a}{c}&0\\ 0&\dfrac{a}{c}(a_{3}b_{5}-a_{5}b_{3})^{2}\end{pmatrix}$
By (2) of Lemma 4.1, we know that $a_{3}b_{5}-a_{5}b_{3}\neq 0$, so the determinant of the matrix has the same sign as $\dfrac{a^{2}}{c^{2}}>0$. We conclude that $Q$ is an elliptic singularity (in the Morse sense). ∎
5. The transition point $P$
In this section, we study the transition point and the transition times.
@@ 5. The transition point $P$
Let $r_{1},\ldots,r_{5}$ be positive numbers such that
  • $\frac{1}{2}r_{5}\leq r_{4}<r_{5}<r_{3}$,
  • $|r_{1}+r_{2}-r_{3}|<r_{5}-r_{4}$, and
  • $|r_{3}-r_{4}-r_{5}|<|r_{1}-r_{2}|<r_{3}+r_{4}-r_{5}$.
Let $j=r_{3}+r_{4}-r_{5}$, and
$t^{\pm}=\frac{\,r_{3}r_{4}+r_{3}r_{5}+2r_{4}^{2}-3r_{4}r_{5}+r_{5}^{2}\,\pm\,2\sqrt{r_{3}r_{4}r_{5}j}}{\,(r_{3}+r_{5})^{2}+4r_{4}j\,}.$
Define $P\in\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$ represented by the planar pentagon $\rho=(\rho_{1},\rho_{2},\rho_{3},\rho_{4},\rho_{5})$, where
$\rho_{1}=\left(\frac{r_{1}^{2}-r^{2}_{2}+j^{2}}{2j},\frac{\sqrt{\big((r_{1}+j)^{2}-r_{2}^{2}\big)\big(r_{2}^{2}-(r_{1}-j)^{2}\big)}}{2j},0\right),$
$\rho_{2}=\left(\frac{r_{2}^{2}-r_{1}^{2}+j^{2}}{2j},-\frac{\sqrt{\big((r_{1}+j)^{2}-r_{2}^{2}\big)\big(r_{2}^{2}-(r_{1}-j)^{2}\big)}}{2j},0\right),$
$\rho_{3}=(-r_{3},0,0),\rho_{4}=(-r_{4},0,0),\rho_{5}=(r_{5},0,0).$
Then the following holds:
  • for $t\in[0,1]$, $P$ is a rank zero singular point of $(J,H_{t}):=(\ell_{12},t\ell_{34}^{2}+(1-t)\ell_{45}^{2})$,
  • for $t\in[0,t^{-})\cup(t^{+},1]$, $P$ is of elliptic-elliptic type, and
  • for $t\in(t^{-},t^{+})$, $P$ is of focus-focus type.
@@
We first notice that $\ell_{12}(P)=r_{3}+r_{4}-r_{5}$ and $\ell_{34}(P)=r_{3}+r_{4}$, so by Lemma 2.5, $P$ is a rank zero singular point of $(J,H_{t})$ for all $t\in[0,1]$. To check the type of singularity, we need to compute the Hessian matrix of $J,H_{t}$, and the matrix of the symplectic form under a basis. Since $\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$ is a sub-quotient of $S^{2}_{r_{1}}\times S^{2}_{r_{2}}\times S^{2}_{r_{3}}\times S^{2}_{r_{4}}\times S^{2}_{r_{5}}$, we can use the standard coordinates on $({\mathbb{R}}^{3})^{5}$ to induce the coordinates on $\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$. Since the $x$-coordinate of each edge of $\rho$ is nonzero, $(y_{1},z_{1},y_{2},z_{2},y_{3},z_{3},y_{4},z_{4},y_{5},z_{5})$ gives a coordinate chart on the product of spheres. Notice that $T_{P}\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})\cong T_{\rho}\widetilde{\mathcal{P}}(r_{1},r_{2},r_{3},r_{4},r_{5})/V_{\rho}$, where $V_{\rho}$ is the space of infinitesimal vector fields generated by the $SO(3)$-action, so it suffices to find vectors in $T_{\rho}\widetilde{\mathcal{P}}(r_{1},r_{2},r_{3},r_{4},r_{5})$ that descend to a basis of $T_{P}\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$.
@@
Define $W_{1}=\frac{\partial}{\partial z_{1}}-\frac{\partial}{\partial z_{3}}$, $W_{2}=\frac{\partial}{\partial z_{1}}-\frac{\partial}{\partial z_{4}}$, $W_{3}=\frac{\partial}{\partial y_{3}}-\frac{\partial}{\partial y_{4}}$, $W_{4}=\frac{\partial}{\partial y_{4}}-\frac{\partial}{\partial y_{5}}$. Notice that $T_{\rho}\widetilde{\mathcal{P}}(r_{1},r_{2},r_{3},r_{4},r_{5})$ is the kernel of $d\mu_{\rho}$, where $\mu:T_{\rho}\widetilde{\mathcal{P}}(r_{1},r_{2},r_{3},r_{4},r_{5})\to{\mathbb{R}}^{3}$ is the moment map of the $SO(3)$-action, so $W_{i}\in T_{\rho}\widetilde{\mathcal{P}}(r_{1},r_{2},r_{3},r_{4},r_{5})$ for $i=1,2,3,4$. It is straightforward to verify that $W_{1},W_{2},W_{3},W_{4}$ span a four dimensional vector space that intersects $V_{\rho}$ trivially, so $W_{1}],[W_{2}],[W_{3}],[W_{4}]$ form a basis of $T_{P}\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$. Using a similar calculation as in the proof of Proposition 4.2, the Hessian of $H_{0}$, $H_{1}$, and $J$ under this basis is given by
$\operatorname{Hess}(H_{0})_{P}=\begin{pmatrix}0&0&0&0\\ 0&\frac{r_{5}}{r_{4}}&0&0\\ 0&0&\frac{r_{5}}{r_{4}}&1-\frac{r_{5}}{r_{4}}\\ 0&0&1-\frac{r_{5}}{r_{4}}&\frac{r_{5}}{r_{4}}+\frac{r_{4}}{r_{5}}-2\end{pmatrix},\text{Hess}(H_{1})_{P}=\begin{pmatrix}-\frac{r_{4}}{r_{3}}&1&0&0\\ 1&-\frac{r_{4}}{r_{3}}&0&0\\ 0&0&-2-\frac{r_{4}}{r_{3}}-\frac{r_{3}}{r_{4}}&1+\frac{r_{3}}{r_{4}}\\ 0&0&1+\frac{r_{3}}{r_{4}}&-\frac{r_{3}}{r_{4}}\\ \end{pmatrix}$.
$\text{Hess}(J)_{P}=\begin{pmatrix}1-\frac{j}{r_{3}}&1&0&0\\ 1&1-\frac{j}{r_{4}}&0&0\\ 0&0&-j(\frac{1}{r_{3}}+\frac{1}{r_{4}})&\frac{j}{r_{4}}\\ 0&0&\frac{j}{r_{4}}&j(\frac{1}{r_{5}}-\frac{1}{r_{4}})\\ \end{pmatrix},\Omega^{-1}_{P}=\begin{pmatrix}0&0&-r_{3}&-r_{3}\\ 0&0&0&-r_{4}\\ r_{3}&0&0&0\\ r_{3}&r_{4}&0&0\\ \end{pmatrix}$.
Consider
$A_{1,1}:=\Omega_{P}^{-1}\big(\text{Hess}(J)_{P}+\text{Hess}(H_{t})_{P}\big)$
$=\begin{pmatrix}0&0&r_{5}-r_{4}-(2r_{3}+r_{4})t&\frac{r_{3}}{r_{5}}\big((r_{3}+2r_{4}-2r_{5})+(2r_{5}-r_{4})t\big)\\ 0&0&r_{3}+2r_{4}-2r_{5}+(r_{3}+r_{5})t&\frac{2(r_{4}-r_{5})^{2}+r_{3}r_{4}-r_{3}r_{5}-((r_{4}-r_{5})^{2}+r_{3}r_{5})t}{r_{5}}\\ r_{4}-r_{5}+r_{4}t&-r_{3}-r_{3}t&0&0\\ -r_{5}&-2r_{5}+r_{5}t&0&0\end{pmatrix}$
The quadratic polynomial corresponding to its characteristic polynomial is given by $\chi_{1,1}(X)=X^{2}+A(t)\,X+B(t)$, where
$A(t)$ $=[(r_{3}+r_{5})^{2}+2r_{4}j]t^{2}+2\!\left[r_{3}j+(r_{5}-r_{4})(r_{4}-2r_{5})\right]t+2r_{3}\,(r_{3}+3r_{4}-3r_{5})+5(r_{4}-r_{5})^{2}$
$B(t)$ $=j^{2}\,\bigl(-r_{4}t^{2}+r_{3}t+r_{4}t+r_{5}t+r_{3}+2r_{4}-2r_{5}\bigr)^{2}.$
@@
By Definition 3.1, $P$ is of elliptic-elliptic type if $\chi_{1,1}(X)$ has two distinct negative roots, i.e. $A(t)>0,B(t)>0,\;A(t)^{2}>4B(t)$; and $P$ is of focus-focus type if $\chi_{1,1}(X)$ has no purely real roots, i.e. $A(t)^{2}<4B(t)$.
We first compare $A(t)^{2}$ with $4B(t)$.
$A(t)^{2}-4B(t)=\bigl((r_{3}+r_{5})t+2r_{3}+3r_{4}-3r_{5}\bigr)^{2}\,\bigl(at^{2}+bt+c\bigr),$
@@
where $a=(r_{3}+r_{5})^{2}+4r_{4}j,b=-2\bigl(r_{3}r_{4}+r_{3}r_{5}+2r_{4}^{2}-3r_{4}r_{5}+r_{5}^{2}\bigr),c=(r_{4}-r_{5})^{2}$. Since $r_{3}>r_{5}$ and $r_{4}\geq\frac{1}{2}r_{5}$, we have $(r_{3}+r_{5})t+2r_{3}+3r_{4}-3r_{5}=(r_{3}+r_{5})t+2(r_{3}-r_{5})+3r_{4}-r_{5}>0$. Hence, $A(t)^{2}-4B(t)$ has the same sign as $f(t):=at^{2}+bt+c$. Since the discriminant $\Delta$ of $f(t)$ is $\Delta=b^{2}-4ac=16\,r_{3}r_{4}r_{5}j>0$ and $a>0$, $f(t)$ has two distinct roots; moreover, by the root formula, the two roots are $t^{\pm}$. We conclude that
  • when $t\in(-\infty,t^{-})\cup(t^{+},\infty)$, we have $A(t)^{2}>4B(t)$, so $\chi_{1,1}(X)$ has two distinct real roots.
@@
  • when $t\in(t^{-},t^{+})$, we have $A(t)^{2}<4B(t)$, so $\chi_{1,1}(X)$ has no real roots, and it follows that $P$ is of focus-focus type.
Moreover, since $r_{3}>r_{5}>r_{4}$, we have $b=-2\big(r_{4}(r_{3}-r_{5})+r_{5}(r_{3}-r_{4})+r_{5}(r_{5}-r_{4})+2r_{4}^{2}\big)<0$ and $a+b=r_{3}^{2}-r_{5}^{2}+2r_{3}r_{4}+2r_{4}r_{5}>0$. Therefore, we know that $f(0)=c>0,f^{\prime}(0)=b<0,f(1)=a+b+c>0,f^{\prime}(1)=2a+b>0$, and hence the two roots $t^{\pm}$ of $f(t)$ must both be in $(0,1)$.

### v2

@@ 4. Morse Functions on quadrilateral spaces
$W_{1}(\widetilde{W}_{i}H_{t})(\rho)=-tg_{3}^{i}(\rho)\frac{b_{4}}{b_{3}}+(1-t)g_{5}^{i}(\rho)\frac{b_{4}}{b_{5}},$
which evaluates to $\frac{b_{4}}{b_{3}}t-\frac{b_{4}}{b_{5}}(1-t)$ when $i=1$ and evaluates to 0 when $i=2$. Similarly, one can compute
$W_{2}(\widetilde{W}_{2}H_{t})(\rho)=\frac{a(a_{3}b_{5}-a_{5}b_{3})^{2}}{c}-t\frac{a_{3}^{2}+b_{3}^{2}}{b_{3}}b_{4}b_{5}^{2}-(1-t)\frac{a_{5}^{2}+b_{5}^{2}}{b_{5}}b_{4}b_{3}^{2}.$
Thus, the matrix representation of the Hessian of $H_{t}$ at $m$ under the basis $W_{1}],[W_{2}]$ is
$\begin{pmatrix}-t\frac{b_{4}}{b_{3}}-(1-t)\frac{b_{4}}{b_{5}}&0\\ 0&\frac{a}{c}(a_{3}b_{5}-a_{5}b_{3})^{2}-\frac{t}{b_{3}}(a_{3}^{2}+b_{3}^{2})b_{4}b_{5}^{2}-\frac{1-t}{b_{5}}(a_{5}^{2}+b_{5}^{2})b_{3}^{2}b_{4}\end{pmatrix}$.
We show that this matrix has a positive determinant. Set $X=a_{4}b_{3}-a_{3}b_{4},\ Y=a_{4}b_{5}-a_{5}b_{4}$, then $X+Y=a_{4}(b_{3}+b_{5})-(a_{3}+a_{5})b_{4}=-a_{4}b_{4}+(c+a_{4})b_{4}=cb_{4}$. Since $c\neq 0$, we can write $b_{4}=\frac{X+Y}{c}$. By Equation (*4.1), $ab_{3}=tX,ab_{5}=(1-t)Y$, so we have $X=\frac{ab_{3}}{t},Y=\frac{ab_{5}}{1-t}$. Hence,
@@
Let $K:=\frac{Y}{X}=\frac{tb_{5}}{(1-t)b_{3}}$, then $A=-\frac{a}{cK}(K+1)^{2}$. Similarly, one can compute
@@
By item (2) of Lemma 4.1, $K\neq-1$, so both $A$ and $B$ are nonzero. Hence, $\det(\operatorname{Hess}(H_{t})_{m})=AB=\dfrac{a^{2}}{c^{2}K^{2}}(K+1)^{2}\left[(Ka_{3}b_{5}+a_{5}b_{3})^{2}+(K+1)^{2}b_{3}^{2}b_{5}^{2}\right]>0$. We conclude that $m$ is an elliptic singularity (in the Morse sense).
Case 2: when $b_{4}=0$, then $a_{4}\neq 0$, so we choose the chart with coordinates $(y_{0},z_{0},x_{3},z_{3},y_{4},z_{4},x_{5},z_{5})$. Notice that by (1) of Lemma 4.1, this could only happen when $t=\dfrac{1}{2}$. A similar computation shows that under the basis $W_{1}=\frac{\partial}{\partial z_{0}}-\frac{\partial}{\partial z_{3}}$ and $W_{2}=(a_{5}b_{3}-a_{3}b_{5})\frac{\partial}{\partial y_{0}}-b_{3}b_{5}\frac{\partial}{\partial x_{3}}+b_{3}b_{5}\frac{\partial}{\partial x_{5}}$, the matrix representation of the Hessian of $H_{t}$ is
$\begin{pmatrix}\dfrac{a}{c}&0\\ 0&\dfrac{a}{c}(a_{3}b_{5}-a_{5}b_{3})^{2}\end{pmatrix}$.
By (2) of Lemma 4.1, we know that $a_{3}b_{5}-a_{5}b_{3}\neq 0$, so the determinant of the matrix has the same sign as $\dfrac{a^{2}}{c^{2}}>0$. We conclude that $m$ is an elliptic singularity (in the Morse sense). ∎
5. The transition points
In this section, we study the transition points and the transition times.
@@ 5. The transition points
Let $r_{1},\ldots,r_{5}$ be positive numbers such that $M:=\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$ is a smooth manifold. Let $P,Q,j_{P},j_{Q},t_{P}^{\pm},t_{Q}^{\pm}$ be as in Section 1. Consider the family $F_{t}=(J,H_{t}):=(\ell_{12},t\ell_{34}^{2}+(1-t)\ell_{45}^{2})$ on $M$, where $t\in[0,1]$.
If the inequalities in case (P) in Theorem 1.1 are satisfied, then
  • (1)
    $P$ is a rank zero singular point of $F_{t}$ which is of elliptic-elliptic type for $t\in[0,t_{P}^{-})\cup(t_{P}^{+},1]$ and of focus-focus type for $t\in(t_{P}^{-},t_{P}^{+})$.
If the inequalities in case (Q) in Theorem 1.1 are satisfied, then
  • (2)
    $Q$ is a rank zero singular point of $F_{t}$ which is of elliptic-elliptic type for $t\in[0,t_{Q}^{-})\cup(t_{Q}^{+},1]$ and of focus-focus type for $t\in(t_{Q}^{-},t_{Q}^{+})$.
If the inequalities in case (P+Q) in Theorem 1.1 are satisfied, then both (1) and (2) hold.
@@
The proofs of all three cases are nearly identical, so we will only prove the result in (1) and discuss how to derive the other cases in the last paragraph of this proof. Assume that the inequalities in case (P) in Theorem 1.1 are satisfied. Then $r_{5}>r_{4}$ and $j_{P}>0$.
@@
By Lemma 2.5, $P$ is a singular point of $J$. By Lemma 2.4, $P$ is also a singular point of $H_{t}$ for all $t\in[0,1]$, so it is a rank zero singularity of $F_{t}$. To check the type of singularity, we need to compute the Hessian matrix of $J$ and $H_{t}$, together with the matrix of the symplectic form with respect to a chosen basis. Since $\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$ is a sub-quotient of $S^{2}_{r_{1}}\times S^{2}_{r_{2}}\times S^{2}_{r_{3}}\times S^{2}_{r_{4}}\times S^{2}_{r_{5}}$, we can use the standard coordinates on $({\mathbb{R}}^{3})^{5}$ to induce the coordinates on $\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$. Since $\rho$ is a representative of $P$ with nonzero $x$-coordinates, $(y_{1},z_{1},y_{2},z_{2},y_{3},z_{3},y_{4},z_{4},y_{5},z_{5})$ gives a coordinate chart near $\rho$ on the product of spheres. Notice that $T_{P}\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})\cong T_{\rho}\widetilde{\mathcal{P}}(r_{1},r_{2},r_{3},r_{4},r_{5})/V_{\rho}$, where $V_{\rho}$ is the space of infinitesimal vector fields generated by the $SO(3)$-action, so it suffices to find vectors in $T_{\rho}\widetilde{\mathcal{P}}(r_{1},r_{2},r_{3},r_{4},r_{5})$ that descend to a basis of $T_{P}\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$.
@@
Define $W_{1}=\frac{\partial}{\partial z_{1}}-\frac{\partial}{\partial z_{3}}$, $W_{2}=\frac{\partial}{\partial z_{1}}-\frac{\partial}{\partial z_{4}}$, $W_{3}=\frac{\partial}{\partial y_{3}}-\frac{\partial}{\partial y_{4}}$, $W_{4}=\frac{\partial}{\partial y_{4}}-\frac{\partial}{\partial y_{5}}$. Notice that $T_{\rho}\widetilde{\mathcal{P}}(r_{1},r_{2},r_{3},r_{4},r_{5})$ is the kernel of $d\mu_{\rho}$, where $\mu:S^{2}_{r_{1}}\times S^{2}_{r_{2}}\times S^{2}_{r_{3}}\times S^{2}_{r_{4}}\times S^{2}_{r_{5}}\to{\mathbb{R}}^{3}$ is the moment map of the $SO(3)$-action, so $W_{i}\in T_{\rho}\widetilde{\mathcal{P}}(r_{1},r_{2},r_{3},r_{4},r_{5})$ for $i=1,2,3,4$. It is straightforward to verify that $W_{1},W_{2},W_{3},W_{4}$ span a 4-dimensional vector space that intersects $V_{\rho}$ trivially, so $W_{1}],[W_{2}],[W_{3}],[W_{4}]$ form a basis of $T_{P}\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$. Using a similar calculation as in the proof of Proposition 4.2, the Hessian of $H_{0}$, $H_{1}$, and $J$, together with the matrix of the symplectic form $\Omega_{P}$, under this basis are given by
$\operatorname{Hess}(H_{0})_{P}=\begin{pmatrix}0&0&0&0\\ 0&\frac{r_{5}}{r_{4}}&0&0\\ 0&0&\frac{r_{5}}{r_{4}}&1-\frac{r_{5}}{r_{4}}\\ 0&0&1-\frac{r_{5}}{r_{4}}&\frac{r_{5}}{r_{4}}+\frac{r_{4}}{r_{5}}-2\end{pmatrix},\ \text{Hess}(H_{1})_{P}=\begin{pmatrix}-\frac{r_{4}}{r_{3}}&1&0&0\\ 1&-\frac{r_{3}}{r_{4}}&0&0\\ 0&0&-2-\frac{r_{4}}{r_{3}}-\frac{r_{3}}{r_{4}}&1+\frac{r_{3}}{r_{4}}\\ 0&0&1+\frac{r_{3}}{r_{4}}&-\frac{r_{3}}{r_{4}}\\ \end{pmatrix}$,
$\text{Hess}(J)_{P}=\begin{pmatrix}1-\frac{j_{P}}{r_{3}}&1&0&0\\ 1&1-\frac{j_{P}}{r_{4}}&0&0\\ 0&0&-j_{P}(\frac{1}{r_{3}}+\frac{1}{r_{4}})&\frac{j_{P}}{r_{4}}\\ 0&0&\frac{j_{P}}{r_{4}}&j_{P}(\frac{1}{r_{5}}-\frac{1}{r_{4}})\\ \end{pmatrix},\ \Omega^{-1}_{P}=\begin{pmatrix}0&0&r_{3}&r_{3}\\ 0&0&0&r_{4}\\ -r_{3}&0&0&0\\ -r_{3}&-r_{4}&0&0\\ \end{pmatrix}$.
Fix $\lambda\in{\mathbb{R}}$ such that $\lambda>\max(\frac{r_{5}-r_{4}}{j_{P}},\frac{r_{3}r_{5}}{j_{P}(j_{P}+r_{4})})$ and consider
$A_{\lambda,1}:=\Omega_{P}^{-1}\big(\lambda\text{Hess}(J)_{P}+\text{Hess}(H_{t})_{P}\big)$
$=\begin{pmatrix}0&0&-(2r_{3}+r_{4})t+r_{3}-j_{P}\lambda&\frac{r_{3}}{r_{5}}((2r_{5}-r_{4})t+r_{4}-r_{5}+j_{P}\lambda)\\ 0&0&(r_{3}+r_{5})t+r_{4}-r_{5}+j_{P}\lambda&\frac{-(r_{3}r_{5}+(r_{4}-r_{5})^{2})t+(r_{4}-r_{5})^{2}+\lambda(r_{4}-r_{5})j_{P}}{r_{5}}\\ r_{4}t+(r_{4}-r_{5})\lambda&-r_{3}(t+\lambda)&0&0\\ -r_{5}\lambda&r_{5}(t-\lambda-1)&0&0\end{pmatrix}$.
The quadratic polynomial corresponding to its characteristic polynomial is given by $\chi_{\lambda,1}(X)=X^{2}+A(t)\,X+B(t)$, where
$A(t)$ $=[(r_{3}+r_{5})^{2}+2r_{4}j_{P}]t^{2}+2\!\left[\lambda(r_{3}+r_{5})j_{P}-r_{3}r_{5}-(r_{4}-r_{5})^{2}\right]t+\lambda^{2}j_{P}^{2}+(\lambda j_{P}+r_{4}-r_{5})^{2},$
$B(t)$ $=j_{P}^{2}\,\bigl(-r_{4}t^{2}+r_{3}\lambda t+r_{4}t+r_{5}\lambda t+j_{P}\lambda^{2}+r_{4}\lambda-r_{5}\lambda\bigr)^{2}.$
@@
By Definition 3.1, $P$ is of elliptic-elliptic type if $\chi_{\lambda,1}(X)$ has two distinct negative roots, i.e. $A(t)>0,B(t)>0,\;A(t)^{2}>4B(t)$; and $P$ is of focus-focus type if $\chi_{\lambda,1}(X)$ has no purely real roots, i.e. $A(t)^{2}<4B(t)$.
@@
We first compare $A(t)^{2}$ with $4B(t)$. $A(t)^{2}-4B(t)=\bigl((r_{3}+r_{5})t+r_{4}-r_{5}+2j_{P}\lambda\bigr)^{2}\,\bigl(at^{2}+bt+c\bigr),$ where $a=(r_{3}+r_{5})^{2}+4r_{4}j_{P},b=-2\bigl(r_{4}j_{P}+r_{3}r_{5}+(r_{4}-r_{5})^{2}\bigr),c=(r_{4}-r_{5})^{2}$. Since $\lambda>\frac{r_{5}-r_{4}}{j_{P}}$, we have $(r_{3}+r_{5})t+r_{4}-r_{5}+2j_{P}\lambda>r_{5}-r_{4}>0$ for all $t\geq 0$. Hence, $A(t)^{2}-4B(t)$ has the same sign as $f(t):=at^{2}+bt+c$ when $t\in[0,1]$. Since the discriminant of $f(t)$ is $b^{2}-4ac=16\,r_{3}r_{4}r_{5}j_{P}>0$ and $a>0$, $f(t)$ has two distinct roots; moreover, by the root formula, the two roots are $t_{P}^{\pm}$. We conclude that
  • when $t\in(-\infty,t_{P}^{-})\cup(t_{P}^{+},\infty)$, we have $A(t)^{2}>4B(t)$, so $\chi_{\lambda,1}(X)$ has two distinct real roots.
@@
  • when $t\in(t_{P}^{-},t_{P}^{+})$, we have $A(t)^{2}<4B(t)$, so $\chi_{\lambda,1}(X)$ has no real roots, and it follows that $P$ is of focus-focus type.
Moreover, since $f(0)=c>0,f^{\prime}(0)=b<0,f(1)=(r_{3}+r_{4})^{2}>0,f^{\prime}(1)=2(r_{3}+r_{4})^{2}+2r_{4}j_{P}+2r_{3}r_{5}>0$, the two roots $t_{P}^{\pm}$ of $f(t)$ must both be in $(0,1)$.